# Incorporating the ChEES Criterion into Sequential Monte Carlo Samplers

### v1

@@ 1 Introduction
The paper proceeds as follows: Section 2 gives an introduction to ChEEs-HMC in the context of MCMC and Section 3 presents SMC and how ChEEs is incorporated as a proposal. Section 4 lays out the experiments with Section 5 providing results and discussion with conclusions and future work given in Section 6.
2 Markov Chain Monte Carlo
@@
2.1 Hamiltonian Monte Carlo
@@
$H(\bm{\theta},\mathbf{p})$ $=-\log\pi(\mathbf{p}|\bm{\theta})-\log p(\bm{\theta})$ (5)
$=-\frac{1}{2}\|\bm{p}\|^{2}-\log p(\bm{\theta}).$ (6)
@@
The leapfrog process is repeated for a certain number of user-specified leapfrog steps $L$ with step size $h$. Pseudocode for the leapfrog algorithm can be found in Algorithm 1. Upon completion of $L$ steps, leapfrog proposes a new sample and momentum, setting $\bm{\theta}^{\prime}=\bm{\theta}_{L}$ and $\mathbf{p}^{\prime}=\mathbf{p}_{L}$, which are accepted according to the criterion
@@
2.2 No-U-Turn Sampler
The most popular adaptive trajectory length selection algorithm is NUTS [16]. At each MCMC iteration, NUTS builds up a binary tree, using the leapfrog algorithm to build up trajectories in both directions alternately while doubling the number of leapfrog steps each time it considers switching direction. The construction of the tree stops when the trajectory makes a "U-turn",
@@ 2.2 No-U-Turn Sampler
2.3 Change in the Estimator of the Expected Square
@@
$\textbf{ChEES}=\frac{1}{4}\mathbb{E}[(||\bm{\theta^{\prime}}-\mathbb{E}[\bm{\theta}]||^{2}-||\bm{\theta}-\mathbb{E}[\bm{\theta}]||^{2})^{2}]$ (12)
This criterion is designed to maximize the change in variance estimation during sampling. Maximizing this reduces the autocorrelations between samples and encourages the sampler to explore the distribution more thoroughly. To aid in this exploration, trajectory lengths are also jittered using a Halton sequence [13]. Gradient descent is used to maximise the ChEES criterion during the warm up period by optimising the trajectory length $L$ hyperparameter and is fixed after this warm up period.
ChEES is often less computationally expensive than NUTS while still being able to adapt the trajectory length effectively and explore the distribution. The benefit of using it in an SMC setting is that we can use multiple samples and adapt the trajectory length in parallel during the warm up phase. Algorithm 20 gives pseudocode for ChEES.
@@
5:     Propose new sample $\bm{\theta}^{(c)^{\prime}}$ and momentum $\mathbf{p}^{(c)^{\prime}}$ using $\text{leapfrog}(\bm{\theta}^{(c)}_{m-1},\mathbf{p}^{(c)^{\prime}},\epsilon,\lceil t_{m}/\epsilon\rceil)$.
@@
[…]
@@
3 Sequential Monte Carlo
@@
At $j=1$, $J$ samples $\forall j=1,\dots,J$ are drawn from a prior distribution $q_{0}(\cdot)$ as follows:
@@
$\mathbf{w}^{j}_{k}=\mathbf{w}^{j}_{k-1}\frac{\pi(\bm{\theta}^{j}_{k})}{\pi(\bm{\theta}^{j}_{k-1})}\frac{L(\bm{\theta}^{j}_{k-1}|\bm{\theta}^{j}_{k})}{q(\bm{\theta}^{j}_{k}|\bm{\theta}^{j}_{k-1})},\ \ \forall i.$ (17)
@@
[…]
@@
        Resample $[\bm{\theta}^{1}_{k}...\bm{\theta}^{J}_{k}]$ with probability $[\tilde{\mathbf{w}}^{1}_{k}...\tilde{\mathbf{w}}^{J}_{k}]$
@@ 3.1 Proposals in SMC
3.2 SMC-ChEES
SMC-ChEES is an extension of the HMC proposals discussed in section 2, as it provides a principled alternative trajectory adaption technique to NUTS. When using this in an SMC context, we can use $N$ random seeds to jitter each trajectory length by a different amount at each iteration, and then use this to compute the acceptance rate $\alpha$ needed for the optimization procedure to maximize ChEES during the warm up phase. Even after warm up, although we fix $L$ we can still use different jitters for each sample to effectively explore the distribution.
3.2.1 Random Number Generators (RNGs)
In [15] the random number sequence $h_{1:M}$ was generated from a 1-dimensional Halton sequence [13]. We explore the usage of the following random and quasi-random number generation schemes to generate a matrix of $N\times K$ random numbers $h_{1:NK}$ for use with ChEES as a proposal in an SMC sampler.
@@ 3.2.1 Random Number Generators (RNGs)
    No jitter: all $N\times K$ numbers are set to 1.
@@
    $h_{nk}=\sum_{d=0}^{\infty}\text{digits}^{k}_{d}(n)k^{-d-1}$ (27)
    where $\text{digits}_{d}(n)$ is ${}^{\text{th}}$ digit of $n$ represented in base-$k$ with the order of the digits reversed.
@@
    1-d Halton: $N\times K$ numbers are taken from (27) with base set to 2.
@@
    Equidistant: $N$ points are created from $h_{n}=n/N$. These $h_{n}$ points are shuffled $K$ times to fill the $h_{nk}$ matrix.
@@
    $h_{nk}=\text{digits}_{1}^{2}(n)\nu_{1}^{k}\oplus\text{digits}_{2}^{2}(n)\nu_{2}^{k}\oplus\dots$ (30)
@@
    1-d Sobol: $N\times K$ numbers are taken from (27) with $k=1$.
4 Experimental Set-up
@@
Table 1: Number of gradient evaluations per sample (smaller is better) and effective sample size per gradient evaluation (larger is better) averaged across iterations.
@@ 4.4 German Credit
$\sigma(x)$ =$\overset{\Delta}{=}\frac{1}{1+e^{-x}}$ (42)
5 Results
@@
[…]
@@
In Fig. 1 the methods alternate for which has the highest $J^{\text{eff}}/\nabla\text{eval}$ across iterations and this is also reflected by Table 1. Our results support previous findings that jittering the trajectory length is helpful in drawing good samples from the posterior distribution [24].
@@
Table 2: Logistic Regression Results for the German Credit Dataset (larger is better).
6 Conclusions
@@
We have also investigated the different RNG methods for jittering outperform no-jittering within the ChEES algorithm with the best overall performance coming from the 1-d Halton and 1-d Sobol methods. We note that the depth has been reached for many methods on the Ill-conditioned Gaussian example which requires more investigation. Despite hitting the maximum number of leapfrog steps, the MSE performs is still on par with NUTS.

### v2

@@ I Introduction
The paper proceeds as follows: Section II gives an introduction to ChEES-HMC in the context of MCMC and Section III presents SMC and how ChEES is incorporated as a proposal. Section IV lays out the experiments with Section V providing results and discussion with conclusions and future work given in Section VI.
II Markov Chain Monte Carlo
@@
II-A Hamiltonian Monte Carlo
@@
$H(\bm{\theta},\mathbf{p})$ $=-\log\pi(\mathbf{p}|\bm{\theta})-\log\pi(\bm{\theta})$ (5)
$=-\frac{1}{2}\|\bm{p}\|^{2}-\log\pi(\bm{\theta}).$ (6)
@@
The leapfrog process is repeated for a certain number of user-specified leapfrog steps $L$ with step size $\epsilon$. Pseudocode for the leapfrog algorithm can be found in Algorithm 1. Upon completion of $L$ steps, leapfrog proposes a new sample and momentum, setting $\bm{\theta}^{\prime}=\bm{\theta}_{L}$ and $\mathbf{p}^{\prime}=\mathbf{p}_{L}$, which are accepted according to the criterion
@@
II-B No-U-Turn Sampler
The most popular adaptive trajectory length selection algorithm is NUTS [15]. At each MCMC iteration, NUTS builds up a binary tree, using the leapfrog algorithm to build up trajectories in both directions alternately while doubling the number of leapfrog steps each time it considers switching direction. The construction of the tree stops when the trajectory makes a ”U-turn”,
@@ II-B No-U-Turn Sampler
II-C Change in the Estimator of the Expected Square
@@
[…]
This criterion promotes exploration by maximizing changes in sample variance, thereby reducing autocorrelation. For each chain $c\in\{1,\dots,C\}$, it is evaluated using proposed and previous states $\bm{\theta}^{(c)^{\prime}}$, $\bm{\theta}_{m-1}^{(c)}$, and their means $\hat{\bm{\theta}}^{\prime}$, $\hat{\bm{\theta}}$. The gradient estimate $\hat{g}^{(c)}$ reflects how trajectory length influences sample dispersion, scaled by the jittered length $l_{m}$ and momentum $\bm{p}^{(c)}$.
To enhance exploration, trajectory lengths are jittered via a Halton sequence $h_{m}$, with $l_{m}=h_{m}L_{m-1}$ [13]. During warm-up, ChEES is maximized using gradient descent to adapt $L_{m}$, based on the weighted average gradient:
$\hat{g}=\frac{\sum_{c}\alpha^{(c)}\hat{g}^{(c)}}{\sum_{c}\alpha^{(c)}}$ (13)
This updates $\log L_{m}$ via Adam, with a moving average $\bar{L}$ maintained. After $m=M_{\text{warmup}}$, $L_{m}$ is fixed to $\bar{L}$.
Compared to NUTS, ChEES is often more efficient while still effectively adapting trajectory lengths. In SMC, multiple chains allow trajectory length to be tuned in parallel during warm-up. Algorithm 2 provides the full procedure.
@@
5:  Propose new sample $\bm{\theta}^{(c)^{\prime}}$ and momentum $\mathbf{p}^{(c)^{\prime}}$ using $\text{leapfrog}(\bm{\theta}^{(c)}_{m-1},\mathbf{p}^{(c)^{\prime}},\epsilon,\lceil l_{m}/\epsilon\rceil)$.
@@
12:        $\begin{aligned} \hat{g}^{(c)}&=l_{m}\Big(\|\bm{\theta}^{(c)^{\prime}}-\hat{\bm{\theta}}^{\prime}\|^{2}\\ &\quad-\|\bm{\theta}^{(c)}_{m-1}-\hat{\bm{\theta}}\|^{2}\Big)(\bm{\theta}^{(c)^{\prime}}-\hat{\bm{\theta}}^{\prime})^{\top}\mathbf{p}^{(c)}\end{aligned}$
@@
III Sequential Monte Carlo
@@
At $k=1$, $J$ samples $j=1,\dots,J$ are drawn from a prior distribution $q_{0}(\cdot)$ as follows:
@@
$\mathbf{w}^{j}_{k}=\mathbf{w}^{j}_{k-1}\frac{\pi(\bm{\theta}^{j}_{k})}{\pi(\bm{\theta}^{j}_{k-1})}\frac{L(\bm{\theta}^{j}_{k-1}|\bm{\theta}^{j}_{k})}{q(\bm{\theta}^{j}_{k}|\bm{\theta}^{j}_{k-1})},\ \ \forall j.$ (18)
@@
$\mathbb{E}_{\pi}\left[f(\bm{\theta}_{k})\right]$ $=\int f(\bm{\theta}_{k})\pi(\bm{\theta}_{k})d\bm{\theta}_{k}$
$=\int f(\bm{\theta}_{k})\frac{\pi(\bm{\theta}_{k})}{q(\bm{\theta}_{k})}q(\bm{\theta}_{k})d\bm{\theta}_{k}\approx\tilde{\mathbf{f}}_{k}$ (21)
@@
   Resample $\bm{\theta}^{1}_{k}...\bm{\theta}^{J}_{k}]$ with probability $\tilde{\mathbf{w}}^{1}_{k}...\tilde{\mathbf{w}}^{J}_{k}]$
@@ III-A Proposals in SMC
III-B SMC-ChEES
SMC-ChEES is an extension of the HMC proposals discussed in section 2, as it provides a principled alternative trajectory adaption technique to NUTS. When using this in an SMC context, we can use $J$ random seeds to jitter each trajectory length by a different amount at each iteration, and then use this to compute the acceptance rate $\alpha$ needed for the optimization procedure to maximize ChEES during the warm up phase. Even after warm up, although we fix $L$ we can still use different jitters for each sample to effectively explore the distribution.
III-B1 Random Number Generators (RNGs)
In [16] the random number sequence $h_{1:M}$ was generated from a 1-dimensional Halton sequence [13]. We explore the usage of the following random and quasi-random number generation schemes to generate a matrix of $J\times K$ random numbers $h_{1:JK}$ for use with ChEES as a proposal in an SMC sampler.
@@ III-B1 Random Number Generators (RNGs)
    No jitter: all $J\times K$ numbers are set to 1.
@@
    $h_{jk}=\sum_{d=0}^{\infty}\text{digits}^{k}_{d}(j)k^{-d-1}$ (28)
    where $\text{digits}_{d}(j)$ is the $^{\text{th}} d$igit of $j$ represented in base-$k$ with the order of the digits reversed.
@@
    1-d Halton: $J\times K$ numbers are taken from (28) with base set to 2.
@@
    Equidistant: $J$ points are created from $h_{j}=j/J$. These $h_{j}$ points are shuffled $K$ times to fill the $h_{jk}$ matrix.
@@
    $h_{jk}=\text{digits}_{1}^{2}(j)\nu_{1}^{k}\oplus\text{digits}_{2}^{2}(j)\nu_{2}^{k}\oplus\dots$ (31)
@@
    1-d Sobol: $J\times K$ numbers are taken from (28) with $k=1$.
IV Experimental Set-up
@@
TABLE I: Number of gradient evaluations per sample (smaller is better) and effective sample size per gradient evaluation (larger is better) averaged across iterations.
@@ IV-D German Credit
$\sigma(x)$ $\overset{\Delta}{=}\frac{1}{1+e^{-x}}$ (43)
V Results
@@
Fig. 1 shows the number of effective samples per gradient evaluation ($J^{\text{eff}}/\nabla\text{eval}$) for each of the 4 experiments. It’s clear to that ChEES has a vastly greater $J^{\text{eff}}/\nabla\text{eval}$ than NUTS across all tasks with every RNG. This is reflected in Table I which shows the average number of gradient evaluations per sample and the average $J^{\text{eff}}/\nabla\text{eval}$ across iterations. For the Gaussian, Ill-conditioned Gaussian and Banana targets, NUTS had approximately 4 to 5 times more average $\nabla\text{eval}$ per particle than the ChEES methods and the same difference can be found in the average $J^{\text{eff}}/\nabla\text{eval}$.
A far larger disparity in the number of $\nabla\text{eval}$ per particle and likewise in the $J^{\text{eff}}/\nabla\text{eval}$ can be seen in the German Credit task where NUTS used about 80 times more evaluations than some of the ChEES methods. In our implementation, the maximum tree depth of NUTS was set to $2^{11}$ meaning that no further leapfrog steps were taken beyond that number, even if the No-U-turn criterion in (11) was not met. NUTS took an average of $1952.19$ ($\nabla\text{eval}/N-1$) leapfrog steps as there is also a gradient calculation undertaken prior to the leapfrog steps. This is far greater than on any other task and the maximum tree depth limit was frequently reached during the sampling process. The ChEES proposals, are less effected by their use on real world data and don’t see the same spike in $\nabla\text{eval}/N$.
@@
In Fig. 1 the methods alternate for which has the highest $J^{\text{eff}}/\nabla\text{eval}$ across iterations and this is also reflected by Table I. Our results support previous findings that jittering the trajectory length is helpful in drawing good samples from the posterior distribution [24].
@@
TABLE II: Logistic Regression Results for the German Credit Dataset (larger is better).
VI Conclusions
@@
We have also investigated the different RNG methods for jittering outperform no-jittering within the ChEES algorithm with the best overall performance coming from the 1-d Halton and 1-d Sobol methods. We note that the depth has been reached for many methods on the Ill-conditioned Gaussian example which requires more investigation. Despite hitting the maximum number of leapfrog steps, the MSE performance is still on par with NUTS.
@@
Acknowledgment
AM and JM were funded by a Research Studentship jointly funded by the EPSRC Centre for Doctoral Training in Distributed Algorithms EP/S023445/1. This work was funded by Dstl in collaboration with the Royal Academy of Engineering via ”Dstl-RAEng Research Chair in Information Fusion” under task RQ0000040616. SM thanks Dstl, UK MOD and the Royal Academy of Engineering for supporting this work. The views and conclusions contained in this paper are of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the UK MOD or the UK Government.